9.The principal value of the argument of (1+ ˆš3)(1−i) is____
\\(-5\\fra{\\pi}{12}\\)
\\(\\fra{5\\pi}{12}\\)
\\(\\fra{-\\pi}{12}\\)
\\(\\fra{\\pi}{12}\\)
1
Expert's answer
2020-05-24T17:55:09-0400
"Arg((1+i\sqrt{3})(1-i))=Arg(1+i\sqrt{3})+Arg(1-i)="

"={\pi\over 3}+(-{\pi\over 4})={\pi\over 12}"